consider the quadratic model h(t)=-16t^2+40t+50 for the height (in feet),h, of an object t seconds after the object has been projected straight up into the air. Find the maximum height attained by the object .how much time does it take to fall back to the ground ?assume that it take the same time for going up and coming down.

Answer by solve_for_x(190)

The equation of the model is the equation of a parabola that opens
downward.

The maximum height will correspond to the vertex of the parabola.

The t-coordinate of the vertex is:

t = -b/2a = -40 / (-2*16) = 1.25

Substituting t = 1.25 into the equation gives a height of:

h(1.25) = -16(1.25)^2 + 40(1.25) + 50 = 75 feet

Solution: The maximum height of the object is 75 feet.